# Asymptotic theory of in-context learning by linear attention

### v2

@@ III.5 Learning transition with increasing pretraining task diversity
((a)) $g_{\text{task}}$ for various $\alpha$, log-log scale
((b)) $g_{\text{task}}$ for various $\alpha$, log-linear scale
Figure 4: Plots of $g_{\text{task}}$ against $\kappa$ for linear transformer model vs dMMSE estimator. Figure (a) has plots on log-log scale highlighting the substantial difference in their rates of convergence towards 0. Dashed and solid lines are theory predictions; dots and squares are numerical simulation. Figure (b) has plots on log-linear scale with the addition of the green $\alpha\to\infty$ curve given by (36) demonstrating the phase transition at $\kappa=1$. Parameters: Linear transformer: $d=100-140$, $\tau=0.2\alpha$ throughout, simulations computed over 20 runs; dMMSE: $d=100$, simulations computed over 1000-5000 runs
@@
First, we need to discuss how to measure task generalization. To motivate our solution, we take a closer look at the ridge estimator ($w_{\text{Bayes}}^{\text{ridge}}$ as in eq. 24), which is the Bayes-optimal estimator given the ICL task structure. We analytically characterize $e_{\text{ICL}}$ and $e_{\text{IDG}}$ errors for this estimator in SI Section SI.6, giving
@@
Simulations of $g_{\text{task}}$ as a function of task diversity, $\kappa=k/d$, are shown in Figure 4 for two inference models: dMMSE estimator, and the linear transformer. The dMMSE estimator is included as a benchmark quantifying the performance of a “pure memorizer”, as it is the Bayes-optimal estimator that assumes task vectors only seen in the training test. We see that, as the task diversity parameter $\kappa$ increases, $g_{\text{task}}$ falls for both estimators but at very different rates.
The $g_{\text{task}}$ for the dMMSE estimator falls very slowly as a function of $\kappa$. This is expected because the explicit form of $w_{\text{Bayes}}^{\text{dMMSE}}$, given in eq. 25, is that of a kernel smoother employing a Gaussian kernel as the weighting function around each task vector. Known results about this class of estimators show the performance of the dMMSE estimator on the ICL task suffers from the “curse of dimensionality” [42, 43, 44], i.e. the required number of training samples $k=d\kappa$ would need to be exponential in $d$ to outperform some given $e_{\text{ICL}}$ error tolerance. Intuitively, the IDG task distribution $\mathcal{P}_{\text{test}}=\mathcal{P}_{\text{train}}=\text{Unif}\{w_{1},\cdots,w_{k}\},w_{i}\sim\mathcal{N}(0,I_{d})$ and the ICL task distribution $\mathcal{P}_{\text{test}}=\mathcal{N}(0,I_{d})$ are indistinguishable (by appropriate measures) only if $k$ is exponentially large in $d$. We present more formal arguments in SI Section SI.6.
@@
for constants $c_{1},c_{2}$ dependent on $\alpha,\tau,\rho$, provided in SI Section SI.6. This sharp fall signals a transition of the model to the task generalization regime. To further understand the nature of this transition and the role of $\kappa$ in the solution learned by the linear attention mechanism, consider the regime where $\tau,\alpha\to\infty$ with $\tau/\alpha\equiv c^{\ast}$ kept fixed. Under this setting, we have
@@
This change in analytical behavior indicates a phase transition at $\kappa=1$. We see this phenomena in both Figure 4(a) and Figure 4(b) where the $g_{\text{task}}$ curve plotted against $\kappa$ becomes nondifferentiable at $\kappa=1$ for $\alpha\to\infty$.
We conclude that the linear transformer model is a much more efficient task generalizer than the dMMSE estimator which is the Bayes-optimal memorizer, as shown in Figure 4. The $1/\kappa$ decay in $g^{\text{LT}}_{\text{task}}$ vs the much slower decay of $g^{\text{dMMSE}}_{\text{task}}$ suggests that the linear transformer quickly learns an inference algorithm that generalizes in-context rather than interpolates between training tasks.
@@ IV Experiments with fully-parameterized nonlinear Transformers
((a)) Error curve in $\tau$: Linear attention, full $K,Q,V$ matrices
((b)) Error curve in $\tau$: Nonlinear model, softmax attention only
((c)) Error curve in $\tau$: Nonlinear model, two blocks of softmax attention and 1-layer MLP
((d)) Interpolation threshold follows predicted quadratic $n\propto d^{2}$ scaling in a range of linear and nonlinear architectures.
@@
Figure 5: Experimental verification in full linear attention (a) and nonlinear models (b), (c), of both scaling definition for $\tau$ and double descent behavior in $n$. 5(a), 5(b), 5(c) show error curves against $\tau$ for various architectures, consistent across token dimension $d=20,40,80$. Some deviations arise near double descent in the linear model (a), possibly due to parameter inversion or instabilities, and for large $\tau$ in the deep nonlinear model (c), possibly due to the training procedure. Double-descent phenomena are confirmed: increasing $n$ will increase error until an interpolation threshold is reached. colored dashed lines indicate experimental interpolation threshold for that architecture and $d$ configuration. 5(d) shows that the location of the interpolation threshold occurs for $n$ proportional to $d^{2}$, as predicted by the linear theory. Dots are experimental interpolation thresholds for various architectures, and dashed lines are best fit curves correspond to fitting $\log(n)=a\log(d)+b$, each with $a\approx 2$ (explicitly, $a_{\text{full linear}}=1.87$, $a_{\text{softmax}}=1.66$, $a_{\text{2 blocks}}=2.13$, $a_{\text{3 blocks}}=2.08$). Interpolation threshold was computed empirically by searching for location in $\tau$ of sharp increase in value and variance of training error at a fixed number of gradient steps. Parameters: $\alpha=1,\kappa=\infty,\rho=0.01$. For Figure 5(a), 5(b), and 5(c): variance shown comes from model trained over different samples of pretraining data; lines show averages over 10 runs and shaded region shows standard deviation.
@@
Given that our theoretical results are derived from a simplified linear attention setting, we aim to determine if these insights can be extended to a nonlinear and deep Transformer models with full $K,Q,V$ parameterization. Even though we would not expect the specific algorithm discussed in Section II.4 to transfer to these nonlinear settings, we will test whether our proposed scalings and main qualitative predictions remain. Specifically, we will investigate the following: (1) whether we have identified the correct scaling regime for non-trivial learning in an in-context learning (ICL) task; (2) whether the full Transformer exhibits a sample-wise double descent, where the location of the peak error scales quadratically with input dimensions as predicted by our theory; and (3) whether the transition from memorization to task generalization occurs.
@@
Figure 6: Plot of $g_{\text{task}}$ against $\kappa$ (loglinear scale) for the nonlinear architecture given in Figure 5(c). This demonstrates consistency of $\kappa$ scaling across increasing dimension choices $d=20,40,80$, as well as a transition in task-generalization measure $g_{\text{task}}$ near $\kappa=1$, familiar from the linear theory. Parameters: $\tau=10,\alpha=1,\rho=0.01$. Variance shown comes from 10 models trained over different samples of pretraining data.
@@
Our experiments [^1] Code to reproduce all experiments is available at https://github.com/Pehlevan-Group/incontext-asymptotics-experiments. are done with a standard Transformer architecture, where each sample context initially takes the form given by (4). The fully-parameterized linear transformer (Figure 5(a)) and the softmax-only transformer (Figure 5(b)) do not use a multilayer perception (MLP) in addition to attention. If MLPs are used (e.g. Figure 5(c) and Figure 5(d)), the architecture consists of blocks with: (1) a single-head softmax self-attention with $K,Q,V\in\mathbb{R}^{(d+1)\times(d+1)}$ matrices, followed by (2) a two-layer dense MLP with GELU activation and hidden layer of size $d+1$ [1]. Residual connections are used between the input tokens (padded from dimension $d$ to $d+1$), the pre-MLP output, and the MLP output. We use a variable number of attention+MLP blocks before returning the final logit corresponding to the $(d+1,\ell+1)$th element in the original embedding structure given by (4). The loss function is the mean squared error (MSE) between the predicted label (the output of the model for a given sample $Z$) and the true value $y_{\ell+1}$. We train the model in an offline setting with $n$ total samples $Z_{1},\cdots,Z_{n}$, divided into 10 batches, using the Adam optimizer [45] with a learning rate $10^{-4}$ until the training error converges, typically requiring 10000 epochs [^2] Note that larger $d$ models are often trained for fewer epochs than smaller $d$ models due to early stopping; that said, whether or not early stopping is used in training does not affect either the alignment of error curves in $d$-scaling nor the qualitative behavior (double descent in $\tau$ and transition in $\kappa$.. The structure of the pretraining and test distributions exactly follows the setup for the ICL and IDG tasks described in Section II.
@@
In Figure 5, we plot the average error for a range of fully-parameterized linear (Figure 5(a)) and nonlinear architectures (Figure 5(b) and Figure 5(c)) on the ICL task against the parameter $\tau=n/d^{2}$, which measures the pretraining sample complexity. Notice that the large-$d$ full $K,Q,V$ parameterized linear architecture studied in Figure 5(a) follows the general trend of the theory derived on a simplified parameterization. Further, in all figures, plots across different values of $d$ roughly collapse, suggesting again that our scaling relations capture the observed behavior.
@@
Across all architectures, we observe a double descent in $\tau$. The peak of this curve occurs at the interpolation threshold, which we identify by tracking the value of $\tau$ at which the training loss becomes non-zero (see figure caption). Our theory predicts that the number of samples $n$ at the peak, as well as the interpolation threshold, should scale with $d^{2}$. Indeed, for the fully-parameterized linear transformer in Figure 5(a), the double descent occurs very close to $\tau=1$ for larger $d$, as predicted by the reduced-parameter linear theory. This general quadratic scaling is summarized in Figure 5(d) across a range of architectures. These findings suggest that nonlinear Transformers exhibit a similar interpolation transition, with the same quadratic scaling in sample size as seen in the simplified model.
@@
Finally, we can investigate the effects of task diversity on $g_{\text{task}}$ in a nonlinear model, shown in Figure 6. This provides empirical evidence towards the choice of $\kappa$ scaling, as errors are consistent over a range of token dimension choices. Further, we recover the memorization-to-task-generalization transition in $\kappa$, as observed before in [20]. In Figure 6, we see the value of $\kappa$ for which $g_{\text{task}}$ appears to converge towards 0 is below 1 (as in (36)), suggesting that the more powerful architecture used here may be more “task-efficient” than the linear architecture in achieving memorization-to-generalization transition.
@@ Acknowledgments
Author contributions
Y.M.L. led formal analysis, along with M.L., J.A.Z-V., A.M., and C.P.   M.L. led computational results and verification. C.P. and Y.M.L supervized the project. C.P. conceptualized the project. All authors contributed to key discussions and manuscript drafting.
@@
SI.1 Notation
Sets, vectors and matrices: For each $n\in\mathbb{N}$, $[n]\coloneqq\left\{1,2,\ldots,n\right\}$. The sphere in $\mathbb{R}^{d}$ with radius $\sqrt{d}$ is expressed as $\mathcal{S}^{d-1}(\sqrt{d})$. For a vector $v\in\mathbb{R}^{d}$, its $\ell_{2}$ norm is denoted by $\lVert v\rVert$. For a matrix $A\in\mathbb{R}^{d\times d}$, $\lVert A\rVert_{\mathsf{op}}$ and $\lVert A\rVert_{\mathsf{F}}$ denote the operator (spectral) norm and the Frobenius norm of $A$, respectively. Additionally, $\lVert A\rVert_{\infty}\coloneqq\max_{i,j\in[n]}\lvert A(i,j)\rvert$ denotes the entry-wise $\ell_{\infty}$ norm. We use $e_{1}$ to denote the first natural basis vector $(1,0,\ldots,0)$, and $I$ is an identity matrix. Their dimensions can be inferred from the context. The trace of $A$ is written as $\operatorname{tr}(A)$.
@@
$\operatorname{vec}(E_{1}E_{2}E_{3})=(E_{1}\otimes E_{3}^{\top})\operatorname{vec}(E_{2}),$ (SI.1.1)
@@
$[M]_{\setminus 0}\in\mathbb{R}^{L\times L}$ (SI.1.2)
@@
$X=\big{(}X^{(d)}(u):d\in\mathbb{N},u\in U^{(d)}\big{)},\quad Y=\big{(}Y^{(d)}(u):d\in\mathbb{N},u\in U^{(d)}\big{)},$
@@
$\sup_{u\in U^{(d)}}\mathbb{P}[X^{(d)}(u)>d^{\varepsilon}Y^{(d)}(u)]\leq d^{-D}$
for sufficiently large $d\geq d_{0}(\varepsilon,D)$. If $X$ is stochastically dominated by $Y$, uniformly in $u$, we use the notation $X\prec Y$. Moreover, if for some family $X$ we have $\lvert X\rvert\prec Y$, we also write $X=\mathcal{O}_{\prec}(Y)$.
@@
$\mathbb{P}\left[\,\lvert X-Y\rvert>d^{-\varepsilon}\right]\leq d^{-D}$ (SI.1.3)
@@
SI.2 Moment Calculations and Generalization Errors

### v3

@@ III.5 Learning transition with increasing pretraining task diversity
Figure 4: Plots of $g_{\text{task}}$ against $\kappa$ for linear transformer model vs dMMSE estimator. (a.) has plots on log-log scale highlighting the substantial difference in their rates of convergence towards 0. Dashed and solid lines are theory predictions; dots and squares are numerical simulation. (b.) has plots on log-linear scale with the addition of the green $\alpha\to\infty$ curve given by (38) demonstrating the phase transition at $\kappa=1$. Parameters: Linear transformer: d=100140d=100-140, $\tau=0.2\alpha$ throughout, simulations computed over 20 runs; dMMSE: $d=100$, simulations computed over 1000-5000 runs
@@
First, we need to discuss how to measure task generalization. To motivate our solution, we take a closer look at the ridge estimator ($w_{\text{Bayes}}^{\text{ridge}}$ as in eq.˜26), which is the Bayes-optimal estimator given the ICL task structure. We analytically characterize $e_{\text{ICL}}$ and $e_{\text{IDG}}$ errors for this estimator in SI Section F, giving
@@
Simulations of $g_{\text{task}}$ as a function of task diversity, $\kappa=k/d$, are shown in Figure˜4 for two inference models: dMMSE estimator, and the linear transformer. The dMMSE estimator is included as a benchmark quantifying the performance of a “pure memorizer”, as it is the Bayes-optimal estimator that assumes task vectors only seen in the training test. We see that, as the task diversity parameter $\kappa$ increases, $g_{\text{task}}$ falls for both estimators but at very different rates.
The $g_{\text{task}}$ for the dMMSE estimator falls very slowly as a function of $\kappa$. This is expected because the explicit form of $w_{\text{Bayes}}^{\text{dMMSE}}$, given in eq.˜27, is that of a kernel smoother employing a Gaussian kernel as the weighting function around each task vector. Known results about this class of estimators show the performance of the dMMSE estimator on the ICL task suffers from the “curse of dimensionality” [42, 43, 44], i.e. the required number of training samples $k=d\kappa$ would need to be exponential in $d$ to outperform some given $e_{\text{ICL}}$ error tolerance. Intuitively, the IDG task distribution $\mathcal{P}_{\text{test}}=\mathcal{P}_{\text{train}}=\text{Unif}\{w_{1},\cdots,w_{k}\},w_{i}\sim\mathcal{N}(0,I_{d})$ and the ICL task distribution $\mathcal{P}_{\text{test}}=\mathcal{N}(0,I_{d})$ are indistinguishable (by appropriate measures) only if $k$ is exponentially large in $d$. We present more formal arguments in SI Section F.
@@
for constants $c_{1},c_{2}$ dependent on $\alpha,\tau,\rho$, provided in SI Section F. This sharp fall signals a transition of the model to the task generalization regime. To further understand the nature of this transition and the role of $\kappa$ in the solution learned by the linear attention mechanism, consider the regime where $\tau,\alpha\to\infty$ with $\alpha/\tau\equiv c^{\ast}$ kept fixed. Under this setting, we have
@@
This change in analytical behavior indicates a phase transition at $\kappa=1$. We see this phenomena in both Figures Figure˜4(a.) and Figure˜4(b.) where the $g_{\text{task}}$ curve plotted against $\kappa$ becomes nondifferentiable at $\kappa=1$ for $\alpha\to\infty$.
We conclude that the linear transformer model is a much more efficient task generalizer than the dMMSE estimator which is the Bayes-optimal memorizer, as shown in Figure˜4. The $1/\kappa$ decay in $g^{\text{LT}}_{\text{task}}$ vs the much slower decay of $g^{\text{dMMSE}}_{\text{task}}$ suggests that the linear transformer quickly learns an inference algorithm that generalizes in-context rather than interpolates between training tasks.
@@ IV Experiments with fully-parameterized nonlinear Transformers
Figure 5: Experimental verification in full linear attention (a.) (full $K,Q,V$ matrices) and nonlinear models (b.) (one block softmax attention only), (c.) (two blocks of softmax attention and 1-layer MLP), of both scaling definition for $\tau$ and double descent behavior in $n$. (a.), (b.), (c.) show error curves against $\tau$ for various architectures, consistent across token dimension $d=20,40,80$. Some deviations arise near double descent in the linear model ($a$), possibly due to parameter inversion or instabilities, and for large $\tau$ in the deep nonlinear model (c.), possibly due to the training procedure. Double-descent phenomena are confirmed: increasing $n$ will increase error until an interpolation threshold is reached. colored dashed lines indicate experimental interpolation threshold for that architecture and $d$ configuration. (d.) shows that the location of the interpolation threshold occurs for $n$ proportional to $d^{2}$, as predicted by the linear theory. Dots are experimental interpolation thresholds for various architectures, and dashed lines are best fit curves correspond to fitting $\log(n)=a\log(d)+b$, each with $a\approx 2$ (explicitly, $a_{\text{full linear}}=1.87$, $a_{\text{softmax}}=1.66$, $a_{\text{2 blocks}}=2.13$, $a_{\text{3 blocks}}=2.08$). Interpolation threshold was computed empirically by searching for location in $\tau$ of sharp increase in value and variance of training error at a fixed number of gradient steps. Parameters: $\alpha=1,\kappa=\infty,\rho=0.01$. For (a.), (b.), and (c.): variance shown comes from model trained over different samples of pretraining data; lines show averages over 10 runs and shaded region shows standard deviation.
@@
Given that our theoretical results are derived from a simplified linear attention setting, we aim to determine if these insights can be extended to a nonlinear and deep Transformer models with full $K,Q,V$ parameterization. Even though we would not expect the specific algorithm discussed in Section˜II.4 to transfer to these nonlinear settings, we will test whether our proposed scalings and main qualitative predictions remain. Specifically, we will investigate the following: (1) whether we have identified the correct scaling regime for non-trivial learning in an in-context learning (ICL) task; (2) whether the full Transformer exhibits a sample-wise double descent, where the location of the peak error scales quadratically with input dimensions as predicted by our theory; and (3) whether the transition from memorization to task generalization occurs.
@@
Figure 6: Plot of $g_{\text{task}}$ against $\kappa$ (loglinear scale) for the nonlinear architecture given in Figure fig.˜5(c.). This demonstrates consistency of $\kappa$ scaling across increasing dimension choices $d=20,40,80$, as well as a transition in task-generalization measure $g_{\text{task}}$ near $\kappa=1$, familiar from the linear theory. Parameters: $\tau=10,\alpha=1,\rho=0.01$. Variance shown comes from 10 models trained over different samples of pretraining data.
@@
Our experiments [^1] Code to reproduce all experiments is available at https://github.com/Pehlevan-Group/incontext-asymptotics-experiments. are done with a standard Transformer architecture, where each sample context initially takes the form given by (6). The fully-parameterized linear transformer fig.˜5(a.) and the softmax-only transformer fig.˜5(b.) do not use a multilayer perception (MLP) in addition to attention. If MLPs are used (e.g. fig.˜5(c.) and fig.˜5(d.)), the architecture consists of blocks with: (1) a single-head softmax self-attention with $K,Q,V\in\mathbb{R}^{(d+1)\times(d+1)}$ matrices, followed by (2) a two-layer dense MLP with GELU activation and hidden layer of size $d+1$ [1]. Residual connections are used between the input tokens (padded from dimension $d$ to $d+1$), the pre-MLP output, and the MLP output. We use a variable number of attention+MLP blocks before returning the final logit corresponding to the $(d+1,\ell+1)$th element in the original embedding structure given by (6). The loss function is the mean squared error (MSE) between the predicted label (the output of the model for a given sample $Z$) and the true value $y_{\ell+1}$. We train the model in an offline setting with $n$ total samples $Z_{1},\cdots,Z_{n}$, divided into 10 batches, using the Adam optimizer [45] with a learning rate $10^{-4}$ until the training error converges, typically requiring 10000 epochs [^2] Note that larger $d$ models are often trained for fewer epochs than smaller $d$ models due to early stopping; that said, whether or not early stopping is used in training does not affect either the alignment of error curves in $d$-scaling nor the qualitative behavior (double descent in $\tau$ and transition in $\kappa$.. The structure of the pretraining and test distributions exactly follows the setup for the ICL and IDG tasks described in Section II.
@@
In Figure˜5, we plot the average error for a range of fully-parameterized linear (a.) and nonlinear architectures (b. and c.) on the ICL task against the parameter $\tau=n/d^{2}$, which measures the pretraining sample complexity. Notice that the large-$d$ full $K,Q,V$ parameterized linear architecture studied in (a.) follows the general trend of the theory derived on a simplified parameterization. Further, in all figures, plots across different values of $d$ roughly collapse, suggesting again that our scaling relations capture the observed behavior.
@@
Across all architectures, we observe a double descent in $\tau$. The peak of this curve occurs at the interpolation threshold, which we identify by tracking the value of $\tau$ at which the training loss becomes non-zero (see figure caption). Our theory predicts that the number of samples $n$ at the peak, as well as the interpolation threshold, should scale with $d^{2}$. Indeed, for the fully-parameterized linear transformer in (fig.˜5a.), the double descent occurs very close to $\tau=1$ for larger $d$, as predicted by the reduced-parameter linear theory. This general quadratic scaling is summarized in (fig.˜5d.) across a range of architectures. These findings suggest that nonlinear Transformers exhibit a similar interpolation transition, with the same quadratic scaling in sample size as seen in the simplified model.
@@
Finally, we can investigate the effects of task diversity on $g_{\text{task}}$ in a nonlinear model, shown in Figure˜6. This provides empirical evidence towards the choice of $\kappa$ scaling, as errors are consistent over a range of token dimension choices. Further, we recover the memorization-to-task-generalization transition in $\kappa$, as observed before in [20]. In Figure˜6, we see the value of $\kappa$ for which $g_{\text{task}}$ appears to converge towards 0 is below 1 (as in (38)), suggesting that the more powerful architecture used here may be more “task-efficient” than the linear architecture in achieving memorization-to-generalization transition.
@@ Acknowledgements
Author Contributions. Y.M.L. led formal analysis, along with M.L., J.A.Z-V., A.M., and C.P. M.L. led computational results and verification. C.P. and Y.M.L supervised the project. C.P. conceptualized the project. All authors contributed to key discussions and manuscript drafting.
@@
Appendix A Notation
Sets, vectors and matrices: For each $n\in\mathbb{N}$, $n]\coloneqq\left\{1,2,\ldots,n\right\}$. The sphere in $\mathbb{R}^{d}$ with radius $\sqrt{d}$ is expressed as $\mathcal{S}^{d-1}(\sqrt{d})$. For a vector $v\in\mathbb{R}^{d}$, its $\ell_{2}$ norm is denoted by $\mathinner{\!\left\lVert v\right\rVert}$. For a matrix $A\in\mathbb{R}^{d\times d}$, $\mathinner{\!\left\lVert A\right\rVert}_{\mathsf{op}}$ and $\mathinner{\!\left\lVert A\right\rVert}_{\mathsf{F}}$ denote the operator (spectral) norm and the Frobenius norm of $A$, respectively. Additionally, $\mathinner{\!\left\lVert A\right\rVert}_{\infty}\coloneqq\max_{i,j\in[n]}\mathinner{\!\left\lvert A(i,j)\right\rvert}$ denotes the entry-wise $\ell_{\infty}$ norm. We use $e_{1}$ to denote the first natural basis vector $(1,0,\ldots,0)$, and $I$ is an identity matrix. Their dimensions can be inferred from the context. The trace of $A$ is written as $\operatorname{tr}(A)$.
@@ Appendix A Notation
$\operatorname{vec}(E_{1}E_{2}E_{3})=(E_{1}\otimes E_{3}^{\top})\operatorname{vec}(E_{2}),$ (39)
@@
$M]_{\setminus 0}\in\mathbb{R}^{L\times L}$ (40)
@@
$X=\big(X^{(d)}(u)\mathrel{\mathop{\ordinarycolon}}d\in\mathbb{N},u\in U^{(d)}\big),\quad Y=\big(Y^{(d)}(u)\mathrel{\mathop{\ordinarycolon}}d\in\mathbb{N},u\in U^{(d)}\big),$ (41)
@@
$\sup_{u\in U^{(d)}}\mathbb{P}[X^{(d)}(u)>d^{\varepsilon}Y^{(d)}(u)]\leq d^{-D}$ (42)
for sufficiently large $d\geq d_{0}(\varepsilon,D)$. If $X$ is stochastically dominated by $Y$, uniformly in $u$, we use the notation $X\prec Y$. Moreover, if for some family $X$ we have $\mathinner{\!\left\lvert X\right\rvert}\prec Y$, we also write $X=\mathcal{O}_{\prec}(Y)$.
@@
$\mathbb{P}\left[\,\mathinner{\!\left\lvert X-Y\right\rvert}>d^{-\varepsilon}\right]\leq d^{-D}$ (43)
@@
Appendix B Moment Calculations and Generalization Errors